OpenStudy (barrelracer011):

HELP! ASAP! Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one half x + 5. y = -2x + 2 y = 2x - 14 y = one half x - 8 y = -one half x - 4

OpenStudy (studygurl14):

well parallel lines have equal slopes. so all you do is plug in the slope from the first line (-1/2) into the equation for a line (y= mx + b) so, y = -1/2x + b Then plug in the point for the second line (4, -6) for x and y, solve for b, then plug in the value of b into y = -1/2x + b Understand?

OpenStudy (barrelracer011):

What do you do with the 4,-6 because I got the first part already I'm just confused what to do with the 4,-6

OpenStudy (barrelracer011):

Also what is a recursive function?

OpenStudy (studygurl14):

(x,y) (4,-6) 4 = x -6 = y y = -1/2x + b (-6) = -1/2(4) + b Solve for b, then put the answer you get for b into y = -1/2x + b